Use the Rational Zero Theorem to find all real zeros. +62? - 1 - 30 - 0
Answer #1

Rational Zero Theolem: P is a rational number writen in lowest terms, and of Pla is a zero of fi and a - Polynomial function: graph the polynomial function... 2= -10 ; y = 61013 +66-101+(10)- y = -420 x=-5 i j = (-1)3 +66-57 (+0) 30 y=0 23-3 ; g = (